The perimeter of a rectangle is 140m. The length is 7m more than twice the width. Find the dimensions.
Answer by checkley77(12844)

L=2W+7
P=2L+2W
140=2(2W+7)+2W
140=4W+14+2W
140=6W+14
6W=140-14
6W=126
W=126/6
W=21 m. ANS.
L=2*21+7=42+7=49 m. ans. for the lengtrh.
Proof:140=2*49=2*21
140=98+42
140=140